OpenStudy (anonymous):

Verify the identity sin4(t)=4sin(t)cos^3(t)-4sin^3(t)cos(t)

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

can i show you my work?

OpenStudy (dumbcow):

is that sin^4(t) or sin(4t) ?

OpenStudy (anonymous):

it's sin(4t)

OpenStudy (anonymous):

\[ \sin (4t)=4\sin(t)\cos^3(t)-4\sin^3(t)\cos(t)\]

OpenStudy (anonymous):

do you see my work okay?

OpenStudy (dumbcow):

yes but its for cos^4(t) ?

OpenStudy (anonymous):

oh no i'm sorry i asked the wrong question

OpenStudy (anonymous):

one moment please

OpenStudy (anonymous):

use the power reducing formulas to rewrite each expression as an equivalent expression that does not contain powers of trigonometric functions greater than 1.

OpenStudy (anonymous):

10cos^4x

OpenStudy (anonymous):

can you help me with that?

OpenStudy (dumbcow):

ok 1 error in your work is: \[\cos^2 (2x) = \frac{1 + \cos(4x)}{2}\]

OpenStudy (anonymous):

wait which step ? can you circle it in my work?

OpenStudy (dumbcow):

at this step: \[\cos^4 x = \frac{1 + 2 \cos (2x)+ \cos^2 (2x)}{4}\] where you substituted for cos^2(2x)

OpenStudy (anonymous):

i circled something in the picture

OpenStudy (dumbcow):

yes where you circled is the error

OpenStudy (anonymous):

what did i do wrong? distrubution?

OpenStudy (dumbcow):

im not sure what you did, but the correct substitution is \[\cos^2 (2x) = \frac{1 + \cos(4x)}{2}\]

OpenStudy (anonymous):

i was trying to get the numerator to have the same denomiator so i did 2/2and distrubted that

OpenStudy (dumbcow):

2/2 is just 1 ? common denominator after substitution is 8

OpenStudy (dumbcow):

\[\cos^4(x) = \frac{1+2\cos(2x) + \frac{1+\cos(4x)}{2}}{4} = \frac{2 + 4\cos(2x) + 1+\cos(4x)}{8}\]

OpenStudy (anonymous):

helllo?